Jan types at an average rate of 12 pages per hour. At that rate, how long wil it take jan to type 100 pages?
Answer by edjones(8007)

x/100 = 1/12 proportion
x=100/12
=8 1/3 hrs
=8 hrs 20 min
.
Ed